OpenStudy (anonymous):

Find the critical numbers of the function g(y)=y-1/y^2-y+1 So far I've gotten to -y^2+2y+1/(y^2-y+1)^2=0 My teacher said that the answer is suppose to be (0,2) but when I finish out the problem in not getting that what am I doing wrong?

OpenStudy (alekos):

I take it that your looking for the critical points or stationary points?

OpenStudy (anonymous):

Yes

OpenStudy (alekos):

g'(y) should be -y^2+2y/(y^2-y+1)

OpenStudy (alekos):

so we have g'(y)=0 => y^2-2y = 0 solve for y

OpenStudy (alekos):

do you follow?

OpenStudy (anonymous):

No not really

OpenStudy (alekos):

do you understand how i got g'(y)?

OpenStudy (anonymous):

Yes I get that part but after trying to solve for y after how you've explained it in still not getting the answer

OpenStudy (alekos):

y^2-2y=0 y(y-2)=0 so y=0 and y=2 are the solutions

OpenStudy (alekos):

got that now?

OpenStudy (anonymous):

Ohhhh ok duh thank you

OpenStudy (alekos):

no problem